OpenStudy (anonymous):

Where does the line through (1,0,1) and (4,-2,2) intersect the place x+y+z=6 ?

OpenStudy (amistre64):

create the vector from point to point; and calibrate it to either of the given points to use in the plane equation

OpenStudy (amistre64):

(1, 0,1) -(4,-2,2) -------- <-3, 2, -1> x = 1-3t y = 0+2t z = 1- t plug these in for x, y and z in the plane equation and soplve for t to find the point

OpenStudy (amistre64):



OpenStudy (anonymous):

so t is -2 right ?

OpenStudy (amistre64):

dunno, lets see :)

OpenStudy (amistre64):

1-3t+ 0+2t+1- t = 6 2-3t+2t- t = 6 2-2t = 6 -2t = 4; t = -2, i agree

OpenStudy (amistre64):

x = 1-3t y = 0+2t z = 1- t x = 1-3(-2) = 7 y = 0+2(-2) = -4 z = 1- (-2) = 3 (7,-4,3) should be the point of intersection if we did it right

OpenStudy (anonymous):

so if there is an equation of a plane with three points, for example if a question is like this A(1,-2,4) is parallel to 3i + 5j + 2k, how do i find the vector equation, parametric equation and Cartesian equation?

OpenStudy (amistre64):

is A a point of a vector?

OpenStudy (amistre64):

*of = or

OpenStudy (anonymous):

its a point on the plane

OpenStudy (amistre64):

ok, it a point on the plane; what does "find the vector equation" mean in this context? or are there other points ?

OpenStudy (amistre64):

<3,5,2> is the vector, i see it now

OpenStudy (amistre64):

x = Ax + 3t y = Ay + 5t z = Az + 2t is the vector equation

OpenStudy (amistre64):

i kept seeing that as an equation of the plane .... doh!!

OpenStudy (anonymous):

how do i find the vector, parametric and cartesian equation with three points ? for example three points on a plane have cordinates A(1,-1,0) , B(0,1,1) and C(2,1,-2) referred to an origin. find a vector equation of the plane

OpenStudy (amistre64):

create 2 vectors from the points; cross them to find the normal; and use the normal and a given point to establish the plane

OpenStudy (amistre64):

A(1,-1,0) B(0,1, 1) -C(2,1,-2) -C(2,1,-2) --------------------- <-1,-2, 2> <-2,0,3> these are the vectors produced from Ato C and B to C

OpenStudy (amistre64):

\begin{array}c x&y&z\\ -1&-2&2\\ -2&0&3 \end{array} the cross produces a normal vector that is just the determinante of this matrix

OpenStudy (amistre64):

x(-2(3)-0(2)) -y(-1(3)--2(2)) + z(-1(0)--2(-2)) x(-6) -y(-3+4) + z(-4) <-6,1,-4> i sour normal; might wanna dbl chk it tho

OpenStudy (amistre64):

forgot a negative; <-6,-1,-4>

OpenStudy (amistre64):

we can drop the negatives all the same .... <6,1,4> is just the same vector in the opposite direction 6(x-Px)+1(y-Py)+4(z-Pz) = 0 where (Px,Py,Pz) is any of the given points

OpenStudy (anonymous):

we have to give it in this form \[r= a + \lambda b\]

OpenStudy (amistre64):

ah, that would be parametric i believe. and i aint to confident in those yet

OpenStudy (amistre64):

class is about to start; taking a final in discrete today; good luck :)

OpenStudy (anonymous):

alright, thanks so much for your help ! :)